# Enhancing Vision-Language Model Pre-training with Image-text Pair Pruning Based on Word Frequency

### v1

@@ 1 Introduction
Figure 1: Zero-shot accuracy on ImageNet-1K classification. CLIP is trained on the CC12M dataset [9]. Using our Word-Frequency-based Image-Text Pair Pruning (WFPP), we achieve comparable performance, while using only approximately 77% of the image-text pairs (1.3×1.3$\times$1.3 × speedup). The image encoder is ViT-B-16 [12]. The "ft" is fine-tuning. “Samples seen” refers to the number of samples processed during pre-training.
@@
Our code is available online [^1] https://github.com/MingliangLiang3/WFPP.
@@ 2 Related work
Due to its remarkable zero-shot transferability, the visual-semantic embedding model as a foundational model has received sustained attention from researchers.
In this section, we describe the pre-training methods for vision-language models and discuss related works to accelerate their pre-training.
@@
DeViSE [19] learns visual-semantic embedding from labeled embeddings which are generated from pre-trained skip-gram on 5.7 million documents (5.4 billion words) extracted from wikipedia.org. The semantic knowledge learned from language provides the zero-shot prediction capability of a visual model, which improves performance on unseen data. To enhance visual-semantic embeddings and achieve good zero-shot performance, CLIP and ALIGN scale the data to 400M to learn the better visual-semantic embedding that achieves remarkable zero-shot performance across 27 datasets [1, 4]. These models are pre-trained by contrastive learning, which pushes positive image-text pairs closer to each other and separates negative image-text pairs, aligning the vision and language by acquiring a visual-semantic embedding from the natural language supervision. However, pre-training VLMs on large-scale data are quite expensive, demanding thousands of GPU days [1, 20].
@@
In order to train VLMs effectively, Fast Language-Image Pre-training (FLIP) [2] removes a large portion of image patches to speedup pre-training VLMs. FLIP uses ViT as an image encoder, reducing computation by 2-4$\times$ by removing 50%-75% patches of the image while obtaining better accuracy than the unmasked model [2]. In addition, they randomly masked 50% of the text to pre-train the VLMs, However, this approach does not work well in text encoders, and the performance of zero-shot classification on ImageNet is decreased. Resource-efficient CLIP (RECLIP) [21] employs a smaller version of images for the initial pre-training of CLIP and subsequently fine-tunes the models using larger versions of the images. The pre-training RECLIP with an image size of 64 × 64 reduces compute resource usage by approximately 80% while still outperforming CLIP on image-text retrieval tasks [21]. Subsampling of Frequent Words for Contrastive Language-Image Pre-training (SW-CLIP) [18] proposed a frequency-based word subsampling technique to reduce text length by half for pre-training VLMs, but does not remove image-text pairs.
@@
Metadata-Curated Language-Image Pre-training (MetaCLIP) [14] creates a balanced subset based on the metadata distribution to pre-train VLMs. To balance the training data, MetaCLIP selects image-text pairs from the data pool where the text contains a metadata entry. These metadata entries consist of four components: WordNet synsets, Wiki unigram, Wiki bigram, and Wiki titles. MetaCLIP utilizes a rich metadata source with 500k entries covering a wide range of concepts. However, it only seeks a balance at the level of the entries (metadata categories), which could lead to certain words being under- or over-represented in the sampled training data. In this paper, we prune image-text pairs based on word frequency to create a more balanced subset for pre-training VLMs. Our approach also is easier to implement, as it does not require collecting and filtering thousands of entries or complex curation processes.
@@ 3 Method
$S(t_{j})=\frac{1}{n}\prod_{i=1}^{n}P(w_{i})$ (3)
In this equation, $t_{j}$ represents the $j-th$ text, while $S(t_{j})$ represents both the joint probability of words being discarded and the probability of the text being discarded from the dataset. $n$ is the length of the text, and the maximum value of $n$ is equal to the maximum value of the input text. The advantage of this equation is that it filters out text containing frequent words to create a subset of the dataset with a balanced word distribution.
@@ 4.2 Evaluation
Table 7: Zero-shot accuracy on ImageNet-1K classification. Comparison of unpruned CLIP and pruning with WFPP and the MetaCLIP method using the CC3M dataset. The image encoder is ViT-B-16. “Samples seen” refers to the proportion of the dataset processed during pre-training, with 100% set as 1.00.
@@
Table 8: Zero-shot accuracy on ImageNet-1K classification. We compared data pruning (50% of CC3M) with WFPP and with the MetaCLIP method. The image encoder is ViT-B-16.
@@ 4.4 Impact of Word Frequency Distribution
4.5 Impact of Text Length and Text Length Normalization
$S(t_{j})=\prod_{i=1}^{n}P(w_{i})$ (4)
Figure 2: Word Distribution: Top 300 words in CC12M [9] After pruning 50% of image-text pairs with WFPP (red) and before pruning (blue). Words are ordered by frequency before pruning. WFPP is seen to reduce high-frequency words while retaining low-frequency words.
4.6 Data Analysis
To gain insight into the nature of the impact of WFPP on word-frequency distribution, we visualize word-frequency distribution before and after pruning in Figure 2. The figure reveals that high-frequency words are removed at a higher rate compared to low-frequency words. Notably, the likelihood of retaining a word increases as its frequency decreases. For example, the retention rate for the high-frequency word “person” was 48.9%, while the retention rate for the less frequent word “bag” was 64.4%.
4.7 Vocabulary size analysis
To demonstrate that WFPP maintains the vocabulary diversity, while improving the word-frequency balance, we calculated the vocabulary size before and after applying WFPP. As shown in Table 12, removing 50% of the image-text pairs, the number of vocabulary words with more than 100 occurrences has not decreased substantially. However, the number of words with frequencies between 5 and 100 decreases. Future work should investigate the impact of these words, which might be low, given their low frequencies.
Table 12: Comparison of vocabulary size before and after applying WFPP. WFPP reduced the number of image-text pairs in the CC12M dataset by 50%.
@@ 5 Conclusion and Outlook
Moving forward, using WFPP to prune very large-scale datasets such as LAION-400M [30] would be an interesting direction to explore. As shown in Figure 1, the improvement in zero-shot classification accuracy on ImageNet-1K increases substantially when 60% instead of 50% of CC12M data is retained after pruning. When more than 60% is retained, the rate of improvement falls off. We believe that between 50% to 60%, we are observing the linear scaling law demonstrated in CLIP [1, 20], but after that CC12M offers insufficient image-text pairs containing low-frequency words in order to maintain this rate of improvement. If we are right, it means that starting with a larger data set like LAION-400M, we could improve zero-shot classification accuracy would already exceed 40% with 280 million. The broader implication is that WFPP would maintain and possibly enhance its ability to improve performance as size of the dataset before pruning is increased.
@@ A.1 More Evaluation
Table 13: Zero-shot robustness evaluation:, comparison of zero-shot accuracy performance between CLIP trained on the original data and CLIP trained on data pruned with the MetaCLIP method and with WFPP, on various datasets. The image encoder is ViT-B-16, and the pre-trained dataset is CC3M [8]. The model is fine-tuned for another epoch on the entire dataset.
@@
A.2 Data Examples

### v2

@@ 1 Introduction
Figure 1: Zero-shot accuracy on ImageNet-1K classification. CLIP is trained on the CC12M dataset [9]. Using our Word-Frequency-based Image-Text Pair Pruning (WFPP), we achieve comparable performance, while using only approximately 77% of the image-text pairs (1.3×1.3$\times$1.3 × speedup). The image encoder is ViT-B-16 [12]. The "ft" is fine-tuning. The w/o ft is without fine-tuning. “Samples seen” refers to the number of samples processed during pre-training.
@@
Our code is available online [^1] https://github.com/MingliangLiang3/WFPP.git.
@@ 2 Related work
Due to its remarkable zero-shot transferability, the visual-semantic embedding model as a foundational model has received sustained attention from researchers. In this section, we describe the pre-training methods for vision-language models and discuss related works to accelerate their pre-training.
@@
Vision-Language Models: DeViSE [19] learns visual-semantic embedding from labeled embeddings which are generated from pre-trained skip-gram on 5.7 million documents (5.4 billion words) extracted from wikipedia.org. The semantic knowledge learned from language provides the zero-shot prediction capability of a visual model, which improves performance on unseen data. To enhance visual-semantic embeddings and achieve good zero-shot performance, CLIP and ALIGN scale the data to 400M to learn the better visual-semantic embedding that achieves remarkable zero-shot performance across 27 datasets [1, 4]. These models are pre-trained by contrastive learning, which pushes positive image-text pairs closer to each other and separates negative image-text pairs, aligning the vision and language by acquiring a visual-semantic embedding from the natural language supervision. However, pre-training VLMs on large-scale data are quite expensive, demanding thousands of GPU days [20, 1].
@@
Efficient Language-Image Pre-training for CLIP: Several methods have been proposed to enhance the efficiency of CLIP models. DeCLIP [21] leverages additional self-supervised losses to improve image and text representations which extends CLIP by incorporating intra-modal self-supervision, cross-modal multi-view supervision, and nearest-neighbor supervision. FILIP [22] introduces a cross-modal late interaction module that refines the contrastive objective by using token-wise maximum similarity between visual and textual tokens. UniCLIP [23] unifies inter-domain (image-text) and intra-domain (image-image, text-text) contrastive losses into a single universal embedding space, capturing comprehensive relationships across and within modalities. Knowledge distillation approaches such as TinyCLIP [24] and MoPE-CLIP [25] transfer knowledge from large pre-trained models to smaller ones via cross-modal distillation to reduce the pre-training budget. Additionally, methods like MobileCLIP [26] and ALIP [27] generate synthetic captions using models pre-trained on large datasets (CoCa [28] and OFA_base [29], respectively) to improve CLIP pre-training. Although these proposed pre-training strategies and synthetic caption methods are efficient, they may still exhibit imbalanced word distributions of pre-training data. Applying WFPP to these methods can further improve training efficiency. Moreover, Fast Language-Image Pre-training (FLIP) [2] removes a large portion of image patches to speed up pre-training VLMs. FLIP uses ViT as an image encoder, reducing computation by 2-4$\times$ by removing 50%-75% patches of the image while obtaining better accuracy than the unmasked model [2]. In addition, they randomly masked 50% of the text to pre-train the VLMs, However, this approach does not work well in text encoders, and the performance of zero-shot classification on ImageNet is decreased. Resource-efficient CLIP (RECLIP) [30] employs a smaller version of images for the initial pre-training of CLIP and subsequently fine-tunes the models using larger versions of the images. The pre-training RECLIP with an image size of 64 × 64 reduces compute resource usage by approximately 80% while still outperforming CLIP on image-text retrieval tasks [30]. Subsampling of Frequent Words for Contrastive Language-Image Pre-training (SW-CLIP) [18] proposed a frequency-based word subsampling technique to reduce text length by half for pre-training VLMs, but does not remove image-text pairs.
@@
Data Puning in VLMs: Several data pruning methods have been developed for Natural Language Processing [31, 32] and Computer Vision [33]. However, in this work, we focus on multimedia data pruning. Metadata-Curated Language-Image Pre-training (MetaCLIP) [14] creates a balanced subset based on the metadata distribution to pre-train VLMs. To balance the training data, MetaCLIP selects image-text pairs from the data pool where the text contains a metadata entry. These metadata entries consist of four components: WordNet synsets, Wiki unigram, Wiki bigram, and Wiki titles. MetaCLIP utilizes a rich metadata source with 500k entries covering a wide range of concepts. However, it only seeks a balance at the level of the entries (metadata categories), which could lead to certain words being under or over-represented in the sampled training data. In this paper, we prune image-text pairs based on word frequency to create a more balanced subset for pre-training VLMs. Our approach is also easier to implement, as it does not require collecting and filtering thousands of entries or engaging in complex, time-consuming curation processes.
@@ 3 Method
$S(t_{j})=\frac{1}{n}(1-\prod_{i=1}^{n}P(w_{i}))$ (3)
In this equation, $t_{j}$ represents the $j-th$ text, while $S(t_{j})$ represents both the probability of words being retained and the probability of the text being retained from the dataset. $n$ is the text length, which is also used to normalize the text probability, and the maximum value of $n$ is equal to the maximum value of the input text. The advantage of this equation is that it filters out text containing frequent words to create a subset of the dataset with a balanced word distribution.
@@ 4.2 Evaluation
Table 7: Zero-shot accuracy on ImageNet-1K classification. Comparison of unpruned CLIP and pruning with WFPP and the MetaCLIP method using the CC3M dataset.
@@
Table 8: Zero-shot accuracy on ImageNet-1K classification. We compared WFPP with MetaCLIP, both of them pre-trained on 50% of the CC3M dataset.
@@ 4.4 Impact of Word Frequency Distribution
4.5 Impact of Text Length and Text Length Normalization
$S(t_{j})=1-\prod_{i=1}^{n}P(w_{i})$ (4)
Figure 2: Word Distribution: The top-50 words in CC12M [9] are shown after pruning 50% of image-text pairs using Random (orange) and WFPP (green) methods, and before pruning (black). We then calculate the word percentages for Random and WFPP before and after pruning. Words are ordered by frequency before pruning. The left Y-axis is the number of words and the left Y-axis is the percentage of words which is the number of words before data pruning divided by the number of words after data pruning.
@@ 4.5 Impact of Text Length and Text Length Normalization
Table 12: Vocabulary Size Comparison: Comparison of vocabulary size before and after applying WFPP. WFPP reduced the number of image-text pairs in the CC12M dataset by 50%.
@@
4.6 Data Analysis
To gain insight into the nature of the impact of WFPP on word-frequency distribution, we visualize word-frequency distribution before and after pruning in Figure 2. The figure reveals that high-frequency words are removed at a higher rate compared to low-frequency words for WFPP. In random pruning, each word has about a 50% chance of being pruned. In contrast, under WFPP, the retention probabilities for most of these top 50 high-frequency words are lower than 50%, especially for words like “person”, “illustration”, and “background”, whose retention rates are 34.59%, 22.81%, and 22.31%, respectively. For infrequent words not shown in the figure, such as “connector”, “swords”, and “grille”, the retention rates are higher—84.53%, 55.55%, and 70.35%, respectively.
4.7 Vocabulary size analysis
To demonstrate that WFPP maintains the vocabulary diversity while improving the word-frequency balance, we calculated the vocabulary size before and after applying WFPP. As shown in Table 12, removing 50% of the image-text pairs, the number of vocabulary words with more than 100 occurrences has not decreased substantially. However, the number of words with frequencies between 5 and 100 decreases. Future work should investigate the impact of these words, which might be low, given their low frequencies.
@@ 5 Conclusion and Outlook
Moving forward, using WFPP to prune very large-scale datasets such as LAION-400M [42] would be an interesting direction to explore. As shown in Figure 1, the improvement in zero-shot classification accuracy on ImageNet-1K increases substantially when 60% instead of 50% of CC12M data is retained after pruning. When more than 60% is retained, the rate of improvement falls off. We believe that between 50% to 60%, we are observing the linear scaling law demonstrated in CLIP [1, 20], but after that CC12M offers insufficient image-text pairs containing low-frequency words in order to maintain this rate of improvement. If we are right, it means that starting with a larger data set like LAION-400M, we could improve zero-shot classification accuracy would already exceed 40% with 280 million. The broader implication is that WFPP would maintain and possibly enhance its ability to improve performance as size of the dataset before pruning is increased. It would also be worthwhile to investigate alternative methods for pruning image-text pair data, such as term frequency-inverse document frequency (TF-IDF).
@@ A.1 More Evaluation
Table 13: Zero-shot robustness evaluation: Comparison of zero-shot accuracy performance between CLIP trained on the original data and CLIP trained on data pruned with the MetaCLIP method and with WFPP, on various datasets. The image encoder is ViT-B-16, and the pre-trained dataset is CC3M [8]. The model is fine-tuned for another epoch on the entire dataset.
@@
A.2 Threshold Selection for WFPP
Based on the idea that high-frequency and low-frequency words should be treated differently, we select the threshold using Equation 2. We aim to decrease the sampling probability of high-frequency words while retaining low-frequency words. For words with frequency $f(w_{i})$ less than $t$ (i.e., very rare words), we set $P(w_{i})$ to 1. These very rare words do not affect the probability of the text being discarded. Specifically, for the cc12m dataset, setting $t$ to 1e-6 means words with count less than 206 are assigned $P(w_{i})=1$; setting $t$ to 1e-7 corresponds to words with a count less than 20 being assigned $P(w_{i})=1$; when $t$ is set to 1e-8, all words are considered by Equation 2 (since no words have frequency less than $t$). We also provide zero-shot classification accuracy with three different thresholds: 1e-6, 1e-7, and 1e-8, As shown in Table 16. When we set the threshold $t$ to 1e-6, the zero-shot classification accuracy is lower than when $t$ is 1e-7 or 1e-8, but it still outperforms the random pruning method. The performances for thresholds $1\times 10^{-7}$ and $1\times 10^{-8}$ are similar.
Table 16: Zero-shot classification accuracy on ImageNet-1K. We pre-trained models by sampling image-text pairs sorted according to Equation 3, with sampling rates of 50% based on different thresholds $t$, specifically 1e-6, 1e-7, and 1e-8. The pre-training dataset is CC12M [9], and the image encoder used is ViT-B-16 [12]. “Samples seen” refers to the proportion of the dataset processed during pre-training, with 100% set as 1.00. w/ft and w/o/ft is with and without fine-tuning.
Method Sample Size threshold w/o/ft w/ft Samples seen (w/ft)
CLIP 4.65M (50%) 28.2 30.2 0.53×0.53$\times$0.53 ×
WFPP 4.65M (50%) 1e-6 29.0 30.7 0.53×0.53$\times$0.53 ×
4.65M (50%) 1e-7 29.8 31.3 0.53×0.53$\times$0.53 ×
4.65M (50%) 1e-8 29.7 31.2 0.53×0.53$\times$0.53 ×
A.3 synthetic captions
Some works utilize synthetic image captions to pre-train CLIP [26, 27]. While these synthetic captions are of higher quality than web captions, they do not ensure a balanced word distribution. To address this imbalance, WFPP can be applied to the synthetic captions datasets. As shown in Table 17, we generated synthetic captions for the CC3M dataset using BLIP [43]. We then pruned 50% of the synthetic data using random pruning and WFPP to pre-train CLIP. Our results show that WFPP outperforms the random pruning method by 0.4% in zero-shot classification accuracy on the zero-shot ImageNet-1k classification.
Table 17: Zero-shot classification accuracy on ImageNet-1K. We use BLIP [43] to generate synthetic captions for the CC3M dataset. We then prune 50% of these synthetic captions using random and WFPP methods. The model is pre-trained for 30 epochs and fine-tuned on the original dataset for 1 epoch. The image encoder employed is ViT-B-16 [12]. The threshold value $t$ in Equation 2 is set to $10^{-7}$.
Method Data Sample Size w/o/ft w/ft Samples seen (w/ft)
CLIP synthetic captions 50% 7.7 12.6 0.53×0.53$\times$0.53 ×
WFPP 50% 7.8 13.0 0.53×0.53$\times$0.53 ×
WFPP can effectively complement the methods of Efficient Language-Image Pre-training for CLIP As discussed in the related works, several studies have demonstrated their effectiveness in pre-training CLIP. To show that WFPP can effectively complement these methods, we apply WFPP and random pruning to DeCLIP [21]. As shown in Table 18, WFPP outperforms random data pruning by 11.0% before fine-tuning and by 11.6% after fine-tuning on the entire dataset. Unexpectedly, DeCLIP performs poorly when 50% of the CC3M data is randomly pruned, while WFPP works significantly better than the random method. To verify the reliability of our results, we pre-trained the model twice but consistently obtained similar outcomes. Our preliminary results show that WFPP can effectively complement DeCLIP, and we believe that WFPP can also complement other methods for efficient language-image pre-training of CLIP.
Table 18: Zero-shot classification accuracy on ImageNet-1K. We pre-trained DeCLIP on 50% of the CC3M (WFPP and Random pruning) dataset with 28000 steps and fine-tuned 4000 steps on the whole dataset. The image encoder employed is ViT-B-32 [12]. The threshold value $t$ in Equation 2 is set to $10^{-7}$.
Model Method Sample Size w/o/ft w/ft Samples seen (w/ft)
DeCLIP 100% 14.1 26.6M
Random 50% 2.2 3.0 26.6M
WFPP 50% 13.2 13.8 26.6M
A.4 Word and Token Frequency
We also evaluate the model’s performance through token frequency data pruning. As indicated in Table 19, the performance based on word and token frequency is similar. This is because most words correspond to individual tokens. Although some words are tokenized into two or more tokens, this has a limited impact on overall performance.
Table 19: Zero-shot classification accuracy on ImageNet-1K. We pre-train models by sampling image-text pairs sorted according to Equation 3 at sampling rates 50% based on word frequency and token frequency. The pre-training dataset is CC12M [9], and the image encoder used is ViT-B-16 [12].
Method Sample Size tokenizer w/o/ft w/ft Samples seen (w/ft)
WFPP 4.65M (50%) word 29.8 31.3 0.53×0.53$\times$0.53 ×
4.65M (50%) token 29.8 31.5 0.53×0.53$\times$0.53 ×
A.5 Data Analysis
We analyzed the data categories before and after applying WFPP. According to Table 20, there is not a significant change in the percentage of word categories. However, when this data is combined with the analysis presented in Figure 4, it becomes apparent that WFPP prunes more high-frequency noun words while retaining more low-frequency nouns. WFPP improves word distribution by selectively pruning high-frequency words, thereby achieving a more balanced distribution. In addition, we visualize the word distribution for the first and second part of the data after applying WFPP, as shown in Figure 5. The second part has many more high-frequency words than the first part and the high-frequency words percentage of the second part is substantially higher than 50%. Moreover, we also analyzed human-labeled datasets by applying WFPP, such as COCO [34] and VG [44]. As illustrated in Figure 6 and Figure 7, the patterns observed in these figures are similar to those in the CC12M dataset. By applying the WFPP method, more high-frequency words are pruned from the first part of the dataset, while the second part retains a higher number of these words.
Table 20: We analyzed the percentages and counts resulting from random and frequency-based pruning methods. Using the CC12M dataset, we pruned 50% of the original texts. NN means noun, JJ means adjective, and VB means verb.
Method NN JJ VB OTHER Total
Before sampling 103,469,117 (50.34%) 10,245,828 (4.98%) 10,649,943 (5.18%) 81,351,966 (39.61%) 205,716,854
Random 51,648,996 (50.25%) 5,120,215 (4.98%) 5,319,414 (5.18%) 40,666,145 (39.59%) 102,754,770
WPFF 47,149,193 (50.48%) 4,491,732 (4.81%) 5,153,531 (5.52%) 36,596,727 (39.19%) 93,391,183
Figure 4: Word Distribution: The top-100 words in CC12M [9] are shown after pruning 50% of image-text pairs using Random (orange) and WFPP (green) methods, and before pruning (black). We then calculate the word percentages for Random and WFPP before and after pruning. Words are ordered by frequency before pruning. The left Y-axis is the number of words and the left Y-axis is the percentage of words which is the number of words before data pruning divided by the number of words after data pruning.
Figure 5: Word Distribution: The top 100 words from CC12M [9] are presented in two parts: the first and second 50% of the image-text pairs using the WFPP method. We then calculate the word percentages for WFPP before and after pruning. Words are ordered by frequency before pruning. The left Y-axis is the number of words and the left Y-axis is the percentage of words which is the number of words before data pruning divided by the number of words after data pruning.
Figure 6: COCO Dataset Word Distribution: The top 100 words from COCO [34] dataset are presented in two parts: the first and second 50% of the image-text pairs using the WFPP method. We then calculate the word percentages for WFPP before and after pruning. Words are ordered by frequency before pruning. The left Y-axis is the number of words and the left Y-axis is the percentage of words which is the number of words before data pruning divided by the number of words after data pruning.
Figure 7: VG Dataset Word Distribution: The top 100 words from VG [44] dataset are presented in two parts: the first and second 50% of the image-text pairs using the WFPP method. We then calculate the word percentages for WFPP before and after pruning. Words are ordered by frequency before pruning. The left Y-axis is the number of words and the left Y-axis is the percentage of words which is the number of words before data pruning divided by the number of words after data pruning.
Appendix B Data Examples